# Prediction Sets and Conformal Inference with

### v3

@@ A.1 Proofs for Section 2
Then $[a_{0},b_{0}]\subset L(Y_{1},\lambda)\cup L(Y_{2},\lambda)^{c}$, and $[a_{1},b_{1}]\subset L(Y_{2},\lambda)\cup L(Y_{1},\lambda)^{c}$. And both $Y_{1},Y_{2}$ are observationally equivalent under interval censoring, corresponding to the selection mechanism that $Y_{1}=Y^{L}$ and $Y^{2}=Y^{U}$.
@@
which shrinks nicely to $(a,b)$ (Folland,, 1999), and as a reult for some $k^{*}$, $\delta=\int_{A_{k^{*}}}f_{Y^{L},Y^{U}}>0$. Let $P_{\pi/2}$ be the one defined in Lemma A.1 and $P_{\pi/2}^{\prime}$ be the positive measure constructed with the same method for the density function $f_{Y^{L},Y^{U}}\mathbb{I}\left\{A_{k^{*}}^{c}\right\}$. For any $y\in[a,b]$, and any $\lambda>0$, let $0<\varepsilon<\delta/\lambda$, and define the following probability distribution, for any interval $B\subset\mathbb{R}$,
@@
It is clear that $P(Y\in C\mid X=x)\geq P\left([Y^{L},Y^{U}]\subset C\mid X=x\right)$ for all $P\in\mathcal{P}$, and we only need to show the reverse inequality also holds. Let $g_{x}=\,\mathrm{d}P_{Y^{L},Y^{U}\mid X=x}/\,\mathrm{d}\mu$ be the Radon-Nikodym derivative of $P_{Y^{L},Y^{U}\mid X=x}$ with respect to a dominating measure $\mu$ on $\mathbb{R}^{2}$. Assume without loss of generality, $-\infty=u_{0}<l_{1}<u_{1}<\dots<l_{M}<u_{M}<\infty=l_{M+1}$, the half space $\mathbb{R}^{2,+}=\left\{(y_{l},y_{u}):y_{u}\geq y_{l}\right\}$ is partitioned into the following regions,
@@
$A^{\prime\prime}_{m}=\left\{(y_{l},y_{u}):l_{m}\leq y_{l}\leq u_{m},y>l_{m+1}\right\},$
@@ A.2 Proofs for Section 3
    Using the fact that $\mu([a_{0},a_{1}]\bigtriangleup[b_{0},b_{1}])\leq\left\lvert b_{0}-a_{0}\right\rvert+\left\lvert b_{1}-a_{1}\right\rvert$, we have either $\left\lvert\hat{\tau}_{0}(x)-\tau_{0}(x)\right\rvert>\frac{\varepsilon}{2}$ or $\left\lvert\hat{\tau}_{1}(x)-\tau_{1}(x)\right\rvert>\frac{\varepsilon}{2}$. Since $\hat{\tau}_{1}(x)-\hat{\tau}_{0}(x)-(\tau_{1}(x)-\tau_{0}(x))\leq 2\nu<\frac{\varepsilon}{2}$, $\hat{\tau}_{0}(x)$ and $\hat{\tau}_{1}(x)$ are on the same side of $\tau_{0}(x),\tau_{1}(x)$. As a result, either $[a_{0},a_{1}]=[\hat{\tau}_{0},\hat{\tau}_{1}-2\nu]$ or $\left[\hat{\tau}_{0}+2\nu,\hat{\tau}_{1}\right]$ satisfies $\mu\left([a_{0},a_{1}]\bigtriangleup[\tau_{0}(x),\tau_{1}(x)]\right)>\varepsilon$ and $\left\lvert[a_{0},a_{1}]\right\rvert\leq\tau_{1}(x)-\tau_{0}(x)$, then we have $P([a_{0},a_{1}];x)<1-\alpha-\delta$. It follows that $P\left(\hat{\tau}_{0},\hat{\tau}_{1};x\right)<1-\alpha-\delta+c_{2}(2\nu)^{\gamma_{2}}$ So $\left\lvert P_{n}(\hat{\tau}_{0}(x),\hat{\tau}_{1}(x);x)-P(\hat{\tau}_{0}(x),\hat{\tau}_{1}(x);x)\right\rvert>\frac{\delta}{2}>\kappa$. Hence $E_{3,x}E_{0,x}\subset E^{c}$.
@@ A.3 Proofs for results in Section 4
Let $s_{i}=s(Y^{L}_{i},Y^{U}_{i},X_{i};\hat{C})$, then $[Y^{L}_{n+1},Y^{U}_{n+1}]\subset\tilde{C}(X_{n+1})$ if and only if $s_{n+1}\leq\vartheta_{1-\alpha}$. Since $s_{i}$’s are exchangeable for $i\in\mathcal{I}_{2}$, $P\left(s_{n+1}\leq\vartheta_{1-\alpha}\right)\geq 1-\alpha$ by Lemma 2 in Romano et al., (2019). ∎
@@
Fix an $\varepsilon>0$, on the event $\sup_{x}\mu\left(\hat{C}(x)\bigtriangleup C^{*}(x)\right)<\frac{\varepsilon}{2}$, we know that
$\left\lvert s(y^{l},y^{u},x;\hat{C})-s(y^{l},y^{u},x;C^{*})\right\rvert<\frac{\varepsilon}{2},$
for all $y^{l},y^{u}$ and $x$. Let $s_{[j]},s^{*}_{[j]}$ be the corresponding order statistics for $s_{j}$ and $s^{*}_{j},j\in\mathcal{I}_{2}$. Let $s_{[l]}=s_{[l]}$ when $l$ is an integer and $s_{[l]}=s_{[\lfloor l\rfloor+1]}$ when $l$ is not an integer. We have $s_{[(n_{2}+1)(1-\alpha)]}>s^{*}_{[(n_{2}+1)(1-\alpha)]}-\varepsilon$ provided that $s^{*}_{[(n_{2}+1)(1-\alpha)]}-s^{*}_{[(n_{2}+1)(1-\alpha)-1]}<\frac{\varepsilon}{2}$ which happens with high probability for sufficiently large $n_{2}$. Similarly we have $s_{[(n_{2}+1)(1-\alpha)]}<s^{*}_{[(n_{2}+1)(1-\alpha)]}+\varepsilon$. Thanks to the consistency of sample quantile, we have that $s^{*}_{[(n_{2}+1)(1-\alpha)]}\to_{p}Q_{1-\alpha}(s^{*})$ as $n_{2}\to\infty$. By the design of the score function, $[y^{L},y^{U}]\subset C(x)$ if and only if $s(y^{L},y^{U},x;C(x))\leq 0$. Since $Q_{1-\alpha}(s^{*})=\inf\left\{z:P(s^{*}\leq z)\geq 1-\alpha\right\}$, combined with
$P(s^{*}\leq 0)=P\left([Y^{L},Y^{U}]\subset C^{*}\right)\geq 1-\alpha,$
and for any $z<0$, $P(s^{*}\leq z)<1-\alpha$ by Assumption 3.2, we have $Q_{1-\alpha}(s^{*})=0$. Hence $\vartheta_{1-\alpha}=o_{p}(1)$. The proof for the local version is similar and omitted. ∎
@@ Appendix B Uniform consistency of the kernel smoothing estimator
In this section, we show the following kernel smoothing estimator $P_{n}(C;x)$ is uniformly consistent for the conditional probability $P(C;x)=P([Y^{L},Y^{U}]\subset C\mid X=x)$ under some primitive conditions. For all $C=\bigcup_{m=1}^{M}[a_{m},b_{m}]$ where $a_{m}\leq b_{m}<a_{m+1}$,
@@
Let $f(y^{l},y^{u};(a_{m},b_{m})_{m=1}^{M})=\sum_{m=1}^{M}\mathbb{I}\left\{a_{m}\leq Y_{i}^{L}\leq Y^{U}_{i}\leq b_{m}\right\}$ which belongs to the following class of functions indexed by the endpoints $(a_{m},b_{m})_{m=1}^{M}$,
$\mathcal{F}_{M}=\left\{f(\cdot,\cdot;(a_{m},b_{m})_{m=1}^{M^{\prime}}):a_{m}\leq b_{m}<a_{m+1},M^{\prime}\leq M\right\},$
@@
Similarly define the population version $P_{f}(x)=\operatorname{\mathrm{E}}\left[f(Y^{L},Y^{U};(a_{m},b_{m})_{m=1}^{M})\right]$. It is straightforward to show that $\mathcal{F}_{M}$ is a measurable VC class of functions, with a constant envelop function $F=1$ and hence for any probability measure $Q$ over the support of $(Y^{L},Y^{U})$,
@@
where $C,v$ depends only on the VC characteristics of $\mathcal{F}_{M}$ and $r\geq 1$ (van der Vaart and Wellner,, 2013). Then the next lemma follows from Corollary 1 in Einmahl and Mason, (2000) with only minor modifications under some conditions on the smoothing kernel and bandwidth sequence.
@@
$h\to 0,\quad nh^{d}\to\infty,\quad\frac{\left\lvert\log h\right\rvert}{\log\log n}\to 0,\quad\frac{nh^{d}}{\log n}\to\infty.$
@@
Let $\mathcal{X}^{\prime}$ be a compact subset of $\mathcal{X}$ where $p_{X}$ is continuous and bounded away from 00 on $\mathcal{X}^{\prime}$, and suppose $p_{(X,Y^{L},Y^{U})}$ is continuous on $\mathcal{X}^{\prime}\times\mathbb{R}^{2}$. Then under Assumptions B.1 and B.2, with probability $1$,
@@
$\sup_{x\in\mathcal{X}^{\prime}}\sup_{f\in\mathcal{F}_{M}}\left\lvert h^{-d}\operatorname{\mathrm{E}}\left[f(Y^{L},Y^{U};(a_{m},b_{m})_{m=1}^{M})K\left(\frac{x-X}{h}\right)\right]-P_{f}(x)\right\rvert=O(h^{2}).$
@@
As shown in Hansen, (2008) Theorem 8, we can let $\mathcal{X}^{\prime}=\mathcal{X}_{n}$ be an expanding subsets at a suitable rate, and the uniform convergence rate needs to be multiplied by a factor of $\delta_{n}^{-1}$ where $\delta_{n}=\inf_{x\in\mathcal{X}_{n}}p_{X}(x)$.
Here we have dealt with the case when the density function $p_{X,Y^{L},Y^{U}}$ is smooth. For fixed censoring, suppose the discrete part is supported on a known finite set $\left\{(L_{i},U_{i}):i=1,\dots,K\right\}$, as in the empirical studies where the brackets are designed by the survey, then the discrete part can be estimated at parametric rate, which is negligible compared to the nonparametric rate above.

### v4

@@ 0.9 CP 10 (37,280 - 47,224) (51,494 - 64,217) (50,508 - 59,915) (64,290 - 77,790)
7 Conclusion
The object of interest in this paper is a prediction set for an outcome given a set of covariates when this outcome is censored. Censoring is a common issue in economic and other data. We first characterise the oracle prediction set, which is the smallest set that maintains a given miscoverage level under censoring. The characterisation leads to a feasible estimation strategy based on the observed sample. We then provide consistent estimators for this oracle prediction set based on nonparametric estimation of the conditional distribution of the observed lower and upper brackets given the conditioning variables. We allow for the prediction set to consist of multiple intervals and consider both random and fixed censoring. Furthermore, we use recent results from conformal inference to obtain a conformal prediction set that maintains a finite sample miscoverage property using a set-based score function. We find that our procedures perform well with simulated data. We also apply our procedures to data from the US Census. There is increasing interest in prediction set estimation, and several intriguing research directions are related to this paper. One promising avenue is to consider prediction sets under partial identification beyond interval censoring. While conformal inference offers finite-sample guarantees on coverage, it remains a research challenge to explore the more traditional statistical inference on the estimated prediction set. Additionally, the design of conformity scores in the context of partial identification presents an interesting and important topic for further investigation.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process: In preparation for the manuscript, ChatGPT was used to format the figures and tables in the LaTeX document and to provide spelling and grammar checks. GitHub Copilot and Codex, provided code suggestions and auto-completions for the python codes implementing the simulation and empirical study. Refine.ink is also used for proofreading. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ A.1 Proofs for Section 2
Then $a_{0},b_{0}]\subset L(Y_{1},\lambda)\cup L(Y_{2},\lambda)^{c}$, and $a_{1},b_{1}]\subset L(Y_{2},\lambda)\cup L(Y_{1},\lambda)^{c}$. And both $Y_{1},Y_{2}$ are observationally equivalent under interval censoring, corresponding to the selection mechanism that $Y_{1}=Y^{L}$ and $Y^{2}=Y^{U}$.
@@
which shrinks nicely to $(a,b)$ (Folland, 1999), and as a reult for some $k^{*}$, $\delta=\int_{A_{k^{*}}}f_{Y^{L},Y^{U}}>0$. Let $P_{\pi/2}$ be the one defined in Lemma A.1 and $P_{\pi/2}^{\prime}$ be the positive measure constructed with the same method for the density function $f_{Y^{L},Y^{U}}\mathbb{I}\left\{A_{k^{*}}^{c}\right\}$. For any $y\in[a,b]$, and any $\lambda>0$, let $0<\varepsilon<\delta/\lambda$, and define the following probability distribution, for any interval $B\subset\mathbb{R}$,
@@
It is clear that $P(Y\in C\mid X=x)\geq P\left([Y^{L},Y^{U}]\subset C\mid X=x\right)$ for all $P\in\mathcal{P}$, and we only need to show the reverse inequality also holds. Let $g_{x}=\,\mathrm{d}P_{Y^{L},Y^{U}\mid X=x}/\,\mathrm{d}\mu$ be the Radon-Nikodym derivative of $P_{Y^{L},Y^{U}\mid X=x}$ with respect to a dominating measure $\mu$ on $\mathbb{R}^{2}$. Assume without loss of generality, $\infty=u_{0}<l_{1}<u_{1}<\dots<l_{M}<u_{M}<\infty=l_{M+1}$, the half space $\mathbb{R}^{2,+}=\left\{(y_{l},y_{u}):y_{u}\geq y_{l}\right\}$ is partitioned into the following regions,
@@
$A^{\prime\prime}_{m}=\left\{(y_{l},y_{u}):l_{m}\leq y_{l}\leq u_{m},y_{u}>l_{m+1}\right\},$
@@ A.2 Proofs for Section 3
    Using the fact that $\mu([a_{0},a_{1}]\bigtriangleup[b_{0},b_{1}])\leq\left\lvert b_{0}-a_{0}\right\rvert+\left\lvert b_{1}-a_{1}\right\rvert$, we have either $\left\lvert\hat{\tau}_{0}(x)-\tau_{0}(x)\right\rvert>\frac{\varepsilon}{2}$ or $\left\lvert\hat{\tau}_{1}(x)-\tau_{1}(x)\right\rvert>\frac{\varepsilon}{2}$. Since $\hat{\tau}_{1}(x)-\hat{\tau}_{0}(x)-(\tau_{1}(x)-\tau_{0}(x))\leq 2\nu<\frac{\varepsilon}{2}$, $\hat{\tau}_{0}(x)$ and $\hat{\tau}_{1}(x)$ are on the same side of $\tau_{0}(x),\tau_{1}(x)$. As a result, either $a_{0},a_{1}]=[\hat{\tau}_{0},\hat{\tau}_{1}-2\nu]$ or $\left[\hat{\tau}_{0}+2\nu,\hat{\tau}_{1}\right]$ satisfies $\mu\left([a_{0},a_{1}]\bigtriangleup[\tau_{0}(x),\tau_{1}(x)]\right)>\varepsilon$ and $\left\lvert[a_{0},a_{1}]\right\rvert\leq\tau_{1}(x)-\tau_{0}(x)$, then we have $P([a_{0},a_{1}];x)<1-\alpha-\delta$. It follows that $P\left(\hat{\tau}_{0},\hat{\tau}_{1};x\right)<1-\alpha-\delta+c_{2}(2\nu)^{\gamma_{2}}$ So $\left\lvert P_{n}(\hat{\tau}_{0}(x),\hat{\tau}_{1}(x);x)-P(\hat{\tau}_{0}(x),\hat{\tau}_{1}(x);x)\right\rvert>\frac{\delta}{2}>\kappa$. Hence $E_{3,x}E_{0,x}\subset E^{c}$.
@@ A.3 Proofs for results in Section 4
[…]
@@
Suppose $\hat{C}(x)=\sqcup_{j}[\hat{a}_{j}(x),\hat{b}_{j}(x)]$ and $C^{*}(x)=\sqcup_{m}[a_{m}(x),b_{m}(x)]$. Let $\mathcal{X}_{\varepsilon}=\{x\in\mathcal{X}:\min_{m}(b_{m}(x)-a_{m}(x))>\varepsilon\}$, and $(\varepsilon_{k})_{k\geq 1}$ be a positive sequence $\varepsilon_{k}\to 0$, then $P_{X}(\mathcal{X}_{\varepsilon_{k}}^{c})\to 0$ as $\varepsilon_{k}\to 0$. For fixed $\varepsilon>0$, define the event
$A_{n}(\varepsilon):=\Bigl\{\sup_{x\in\mathcal{X}_{\varepsilon}}\mu\bigl(\hat{C}_{n}(x)\triangle C^{*}(x)\bigr)<\varepsilon/2\Bigr\},$
and we have $P(A_{n}(\varepsilon))\to 0$ by the uniform consistency. Fix $\varepsilon>0$ and $x\in\mathcal{X}_{\varepsilon}$, on the event $\sup_{x\in\mathcal{X}_{\varepsilon}}\mu\left(\hat{C}(x)\bigtriangleup C^{*}(x)\right)<\frac{\varepsilon}{2}$, for any $j$, either $\hat{b}_{j}(x)-\hat{a}_{j}(x)<\varepsilon/2$, or $\hat{a}_{j}(x),\hat{b}_{j}(x)]\cap[a_{m}(x),b_{m}(x)]\neq\varnothing$ for some $m$. Also, for any $a_{m}(x),b_{m}(x)]$, there exists at least one $j$, such that $a_{m}(x),b_{m}(x)]\cap[\hat{a}_{j}(x),\hat{b}_{j}(x)]\neq\varnothing$. That is, each true interval $a_{m}(x),b_{m}(x)]$ intersects at least one estimated interval $\hat{a}_{j}(x),\hat{b}_{j}(x)]$: otherwise $a_{m}(x),b_{m}(x)]\subset C^{*}(x)\setminus\hat{C}_{n}(x)$ and hence $\mu(\hat{C}_{n}(x)\triangle C^{*}(x))\geq b_{m}(x)-a_{m}(x)>\varepsilon$, contradicting $A_{n}(\varepsilon)$. For such an intersecting pair, $(\hat{a}_{j}(x)-a_{m}(x))_{+}\leq\mu(C^{*}(x)\setminus\hat{C}_{n}(x))\leq\mu(\hat{C}_{n}(x)\triangle C^{*}(x))<\varepsilon/2,$ and similarly $(b_{m}(x)-\hat{b}_{j}(x))_{+}<\varepsilon/2$, so that $a_{m}(x),b_{m}(x)]\subset[\hat{a}_{j}(x)-\varepsilon/2,\ \hat{b}_{j}(x)+\varepsilon/2]\subset\mathcal{E}(\hat{C}_{n}(x),\varepsilon/2)$. Here $\mathcal{E}(C,t)=\cup_{m=1}^{M}[a_{m}-t,b_{m}+t]$ for $C=\sqcup_{m=1}^{M}[a_{m},b_{m}]\in\mathcal{C}_{M}$. Taking the union over $m$ yields
$C^{*}(x)\subset\mathcal{E}(\hat{C}_{n}(x),\varepsilon/2),\qquad x\in\mathcal{X}_{\varepsilon},\ \text{on }A_{n}(\varepsilon).$
Consequently, for any $t\in(0,\varepsilon/2)$, if $s^{*}(y^{L},y^{U},x)\leq t$, then $y^{L},y^{U}]\subset\mathcal{E}(C^{*}(x),t)\subset\mathcal{E}(\hat{C}_{n}(x),\varepsilon)$ for all $x\in\mathcal{X}_{\varepsilon}$. So that $P(s(Y^{L},Y^{U},x)\leq\varepsilon\mid X=x)\geq P(s^{*}(Y^{L},Y^{U},x)\leq t\mid X=x)>1-\alpha$ for $x\in\mathcal{X}_{\varepsilon}$.
On the other hand, suppose $s(y^{L},y^{U},x)\leq-\varepsilon$ and $y^{L},y^{U}]\subset\mathcal{E}(\hat{C}(x),-\varepsilon)$. If $\mathcal{E}(\hat{C}(x),-\varepsilon)=\varnothing$, then $P(s(y^{L},y^{U},x)\leq-\varepsilon)=0$, otherwise, $\mu(\mathcal{E}(\hat{C}(x),-\varepsilon))\leq\mu(\hat{C}(x))-2\varepsilon\leq\mu(C^{*})-\varepsilon$ on $A_{n}(\varepsilon)$, and
$\mu\!\left(C^{*}(x)\triangle\mathcal{E}(\hat{C}_{n}(x),-\varepsilon)\right)$ $\geq\mu\!\left(\hat{C}_{n}(x)\triangle\mathcal{E}(\hat{C}_{n}(x),-\varepsilon)\right)-\mu\!\left(\hat{C}_{n}(x)\triangle C^{*}(x)\right)$
$\geq 2\varepsilon-\varepsilon/2\geq\varepsilon.$
By Assumption 3.5, there exists $\delta(\varepsilon)>0$ such that for $P_{X}$-a.e. $x$,
$P\Big([Y^{L},Y^{U}]\subset\mathcal{E}(\hat{C}_{n}(x),-\varepsilon)\ \big|\ X=x\Big)\leq 1-\alpha-\delta(\varepsilon).$
So we have the quantile of of $s(Y^{L},Y^{U},X)$ falls in $\varepsilon,\varepsilon]$ with probability approaching $1$. Since the calibration scores $(s_{j})_{j\in\mathcal{I}_{2}}$ are i.i.d. conditional on $\hat{C}_{n}$, standard sample-quantile consistency implies $\hat{\vartheta}_{1-\alpha}\to 0$ in probability. The proof for the local version is similar and omitted. ∎
@@ Appendix B Uniform consistency of the kernel smoothing estimator
In this section, we show the following kernel smoothing estimator $P_{n}(C;x)$ is uniformly consistent for the conditional probability $P(C;x)=P([Y^{L},Y^{U}]\subset C\mid X=x)$ under some primitive conditions. For all $C=\bigcup_{m=1}^{M}[a_{m},b_{m}]$ where $a_{m}<b_{m}<a_{m+1}$,
@@
Let $f(y^{l},y^{u};(a_{m},b_{m})_{m=1}^{M})=\sum_{m=1}^{M}\mathbb{I}\left\{a_{m}\leq y^{L}<y^{U}\leq b_{m}\right\}$ which belongs to the following class of functions indexed by the endpoints $(a_{m},b_{m})_{m=1}^{M}$,
$\mathcal{F}_{M}=\left\{f(\cdot,\cdot;(a_{m},b_{m})_{m=1}^{M^{\prime}}):a_{m}<b_{m}<a_{m+1},M^{\prime}\leq M\right\},$
@@
Similarly define the population version $P_{f}(x)=\operatorname{\mathrm{E}}\left[f(Y^{L},Y^{U};(a_{m},b_{m})_{m=1}^{M})\mid X=x\right]$. It is straightforward to show that $\mathcal{F}_{M}$ is a measurable VC class of functions, with a constant envelop function $F=1$ and hence for any probability measure $Q$ over the support of $(Y^{L},Y^{U})$,
@@
where $C,v$ depends only on the VC characteristics of $\mathcal{F}_{M}$ and $r\geq 1$ (van der Vaart and Wellner, 2013). Then the next lemma follows from Corollary 1 in Einmahl and Mason (2000) with only minor modifications under some conditions on the smoothing kernel and bandwidth sequence.
@@
$h\to 0,\quad nh^{d}\to\infty,\quad\frac{\left\lvert\log h\right\rvert}{\log\log n}\to\infty,\quad\frac{nh^{d}}{\log n}\to\infty.$
@@
Let $\mathcal{X}^{\prime}$ be a compact subset of $\mathcal{X}$ where $p_{X}$ is continuous and bounded away from 0 on $\mathcal{X}^{\prime}$, and suppose $p_{(X,Y^{L},Y^{U})}$ is continuous on $\mathcal{X}^{\prime}\times\mathbb{R}^{2}$. Then under Assumptions B.1 and B.2, with probability $1$,
@@
$\sup_{x\in\mathcal{X}^{\prime}}\sup_{f\in\mathcal{F}_{M}}\left\lvert h^{-d}\operatorname{\mathrm{E}}\left[f(Y^{L},Y^{U};(a_{m},b_{m})_{m=1}^{M})K\left(\frac{x-X}{h}\right)\right]-p_{X}(x)P_{f}(x)\right\rvert=O(h^{2}).$
@@
As shown in Hansen (2008) Theorem 8, we can let $\mathcal{X}^{\prime}=\mathcal{X}_{n}$ be an expanding subsets at a suitable rate, and the uniform convergence rate needs to be multiplied by a factor of $\delta_{n}^{-1}$ where $\delta_{n}=\inf_{x\in\mathcal{X}_{n}}p_{X}(x)$.